OpenStudy (anonymous):

Cathy spent 4/5 of her money while Josie spent 1/2 of her money. Both of them had the same amount of money left. If Josie had $35 left, how much did Cathy have at first?

OpenStudy (anonymous):

if cathy spent 4/5 of her money, she would be left with 1/5 of her money and if they had the smae amount of money left, cathy would also have $35 left

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 }=$35\]\[\frac{ 5 }{ 5 }= ?\]

OpenStudy (jack1):

cathy spent 4/5 of her money, and has $35 josie spent half of her money, so if josie is left with $ 35, she started with 35 x 2/1 = 70$ so if cathy is left with $35, she started with 35 x 5/1 = 175$

OpenStudy (jack1):

josie: x/2 = 35 therefore x = 70 cathy y/5 = 35 therefore y = 175

OpenStudy (anonymous):

1/2x = 35 x = 35/ (1/2) x = 35 x 2/1 = 70 < -- Josie 1/5x = 35 x = 35/(1/5) x = 35 x 5/1 = 175 <-- cathy